OpenStudy (anonymous):

For a closed rectangular box, with a square base x by x cm and height h cm, find the dimensions giving the maximum volume, given that the surface area is 10 cm^2. write EXACT answers... x=______ h=______ thank you!! where do i start?? :O @Luigi0210 :)

OpenStudy (luigi0210):

@ganeshie8 can help.. I don't understand optimization either

OpenStudy (anonymous):

ahh okay thanks!!

OpenStudy (anonymous):

not sure if @ganeshie8 is there hahaa

OpenStudy (agent0smith):

Write out the SA equation, just add up the area of all the sides, and set it equal to 10. Use that equation to solve for h in terms of x. Volume is area of base times height. Differentiate the volume, set equal to zero, solve for x. Then use x to find h.

OpenStudy (anonymous):

would that be l*w*h ?

OpenStudy (agent0smith):

Yes, but use x and h.

OpenStudy (anonymous):

okay like this? x* h * l ? not too sure what that means haha

OpenStudy (anonymous):

how would the SA equation look?

OpenStudy (agent0smith):

I would draw a diagram, it will make it easier for you to visualize it and find the SA. The dimensions are x by x by h, so there's no need for L.

OpenStudy (anonymous):

okay so it will be x * h = surface area?

OpenStudy (anonymous):

(x*h)^4 ?

OpenStudy (agent0smith):

The base and top has area x*x There's 4 sides with area x*h Add them all up

OpenStudy (anonymous):

(x*h)+(x*h)+(x*h)+(x*h) ?

OpenStudy (agent0smith):

You're missing the base and top.

OpenStudy (anonymous):

so (x*h)+(x*h)+(x*h)+(x*h)+(x*h)+(x*h) ?

OpenStudy (agent0smith):

The base and top have area x*x There's 4 sides with area x*h Add them all up

OpenStudy (agent0smith):

Your SA = 4xh + x^2 = 10 solve it for h then follow the steps in the first post.

OpenStudy (agent0smith):

\[\Large V = x^2 h\]

OpenStudy (anonymous):

so 4xh = 10-x^2 h= 10-x^2 ------- 4x ? @agent0smith

OpenStudy (anonymous):

and then\[V=(x ^{2})(\frac{ 10-x^{2} }{ 4x })\] ?

OpenStudy (anonymous):

and then derivative is this? \[V'=\frac{ 5 }{ 2 }-\frac{ 3x^{2} }{ 4 }\] ?

OpenStudy (anonymous):

then set equal to zero?